OpenStudy (anonymous):

Last month, Steven had $4285.68 that he divided into two accounts. He deposited into a savings account. He deposited 1/3 the rest into his checking account. How much money did Steven deposit into his checking account?

OpenStudy (texaschic101):

there are missing numbers in this problem

OpenStudy (anonymous):

Last month, Steven had $4285.68 that he divided into two accounts. He deposited 1/3 into a savings account. He deposited the rest into his checking account. How much money did Steven deposit into his checking account?

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

4285.68 divided into 2 accounts... If 1/3 was put into his savings account, that means 2/3 was put into his checking account. 2/3 * 4258.68 = 8517.36/3 = 2839.12 (this is in his checking account) check.. 1/3 of 4258.68 (savings account) = 4258.68 * 1/3 = 4258.68/3 = 1419.56 2/3 of 4258.68 (checking account) = 4258.68 * 2/3 = 4258.68/3 = 2839.12 1419.56 + 2839.12 = 4258.68 (correct)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Roberto’s monthly pay is $4349.70. Each month he uses 2/5 of his pay for his household expenses and 1/10 for his car payment. He deposits the rest into his savings account. How much does Roberto deposit into his savings account each month

OpenStudy (texaschic101):

4349.70 (monthly pay) 2/5 * 4349.70 = 8699.40/5 = 1739.88 (household expenses) 1/10 * 4349.70 = 4349.70/10 = 434.97 (car payment) 4349.70 - (1739.88 + 434.97) = amount deposited into savings account 4349.70 - (2174.85) = savings account 2174.85 = amount in savings account